OpenStudy (anonymous):

How do you factor 25x^3 - 15x^2

OpenStudy (anonymous):

okay well you have to factor out common factors do you know how to do that?

OpenStudy (midhun.madhu1987):

Welcome to OpenStudy... 25x^3 - 15x^2 25x^3 = 5 * 5 * x * x * x 15x^2 = 3 * 5 * x * x Now take out the common factors from both terms...

OpenStudy (anonymous):

No, this is all review but I forgot everything in the past year or two.

OpenStudy (anonymous):

Ohhhh okay. Thank you so much.

OpenStudy (midhun.madhu1987):

ur welcome :)

OpenStudy (anonymous):

Wait, so would the answer be just 5x-3?